# Mixed-Integer Programming STALC Stratified Topological Autonomy for Long-Range Coordination

### v1

@@ VI-A Mixed-Integer Programming Approach
$n_{E}$ Number of edges, both directions
@@
$n_{S}$ Number of start locations
$n_{G}$ Number of goal locations
@@
$S$ Set of start locations $s$, $S\subseteq L$
@@
${O}$ Set of overwatch opportunities ${o}$, $(v_{i},e_{j})$ where node $v_{i}$ can overwatch edge $e_{j}$
@@
$n_{s}$ Number of robots at start location $s\in S$
@@
Cost of Traversing $w_{e}$ Base cost to traverse edge $e\in E$
@@
$\alpha_{o}$ Number of robots for full overwatch for ${o}$
$\gamma_{o}$ Reward for overwatch robots over $\alpha_{o}$ for ${o}$
@@
$\breve{C}_{W_{e,t}}$ Cont. 0 $\infty$ Cost of traversing edge $e$
$\breve{C}_{\Omega_{{o},t}}$ Cont. $\infty$- ∞ 0 Cost of overwatch opportunity ${o}$
VI-B Key Formulation Innovations
We apply three core innovations to preserve linearity in our expression of the MIP and significantly improve computation time: our integer decision variable formulation, utilization of indicator variables, and methods for maintaining convexity in our cost functions.
VI-B1 Integer Decision Variable Formulation
Common formulations for multi-agent problems consider all possible states of robots at locations in the graph across all time steps. This is $n_{T}n_{L}^{n_{A}}$ possible states. In a MIP formulation, enumerating these states as binary decision variables would be computationally intractable. Another common formulation is to encode each robot’s location at each time step. This enumerates as $n_{T}n_{L}n_{A}$ total variables which is also highly computationally intensive for large numbers of agents and locations.
We drastically reduce the size of the decision space by instead considering how many robots are at each location at each time step with $n_{T}n_{L}$ integer variables. This still allows us to analyze interactions between robots sharing edges or providing overwatch, but eliminates the exponential growth of the decision space. This formulation obfuscates the paths of each individual robot in the optimization program, though they can easily be assigned afterwards. Additionally, the paths can be ensured to follow the topology of the graph through sequential flow constraints considering the movement of all robots, which we present in the following sections.
With this formulation, we embed the multi-robot interactions directly in the edge weights of our dynamic topological graph. In particular, we consider benefits from moving in formations and as a team to directly scale with the number of robots on an edge; thus our choice of decision variable greatly simplifies these calculations. For interactions through overwatch opportunities, we define costs and constraints for each overwatch opportunity that depend on the number of robots at the overwatch nodes and related edges to reduce the cost of traversing those edges. Fully defining each of these possible interactions when instead enumerating variables for each robot would significantly increase the problem’s complexity. Our choice of decision variables is pivotal to keeping the problem computationally tractable in a complex decision space.
VI-B2 Indicator Variables
To maintain linearity in our problem formulation, we add binary, or indicator, decision variables when necessary using the following form.
Proposition 1.
For $b\in[0,b^{\ulcorner}]$, an indicator variable $\alpha\in\{0,1\}$ can be expressed as follows, for any constant $a\geq 0$.
$\alpha=\begin{cases}1,~{}~{}b>0\\ 0,~{}~{}b=0\end{cases}~{}~{}~{}\Leftrightarrow~{}~{}~{}~{}~{}\min a\alpha~{}\text{s.t.}~{}\alpha\geq\frac{b}{b^{\ulcorner}}$ (10)
We can then use the newly created binary variable $\alpha$ in our cost functions and constraints. As long as $\alpha$ is used in contexts that will be minimized, the value of $\alpha$ will be strict to the piecewise expression in (10). This is due to the formulation; $\alpha$ can only be 00 or $1$, and the constraint follows that $\frac{b}{b^{\ulcorner}}>0$ for $b>0$ and $\frac{b}{b^{\ulcorner}}=0$ for $b=0$. For $\alpha$ to be minimized means that the associated cost function has the lowest value when $\alpha=0$ (compared to $\alpha=1$). Indicator variables of this form can aid in expressing complex equations more efficiently and effectively in our optimization problem.
VI-B3 Convex Cost Functions
We aim to express our costs and constraints convexly, since convex cost functions can significantly improve the computation time of an optimization program. In particular, since we are using MIP, we want to formulate our cost functions to be convex when our integrality constraints are relaxed. For brevity, when we refer to the convexity of a function, we are considering the convexity when the integrality constraints are relaxed.
To produce a convex and positively homogeneous function, we can form a perspective function, as originally presented in [36]. In particular, we utilize the following result from [37], where $\mathbb{R}_{>0}$ is the set of positive real numbers.
Proposition 2.
$\tilde{f}(x,y)=yf\Big{(}\frac{x}{y}\Big{)},\quad x\in\mathbb{R}^{n},~{}y\in\mathbb{R}_{>0}$ (11)
The following example from [38] shows a perspective function in practice.
Example 1.
For $y\geq 0$, the perspective function of the convex function $f(x)=||Ax+b||$ is given by $\tilde{f}(x,y)=||Ax+by||$, where $x\in\mathbb{R}^{n},~{}A\in\mathbb{R}^{n\times n},~{}b\in\mathbb{R}^{n}$. $\tilde{f}$ is jointly convex in $x$ and $y$ by Proposition 2.
In this work, we formulate our cost functions as piecewise functions based on the different regimes of operating alone or with a team. Using perspective functions, we can derive convex piecewise cost functions. This allows us to express these complex functions compactly through a linear cost term with linear constraints as follows.
Proposition 3.
If $f(x)$ is a convex piecewise linear function with constraints $c_{1}(x),\dots,c_{n}(x)$, then the following are equivalent for some auxiliary variable $\breve{f}$.
$\min f(x)\quad\Leftrightarrow\quad\min\breve{f}~{}\text{s.t.}~{}\breve{f}\geq c_{1}(x),\dots,\breve{f}\geq c_{n}(x)$ (12)
VI-C MIP Cost Function
Using the information embedded in our dynamic topological graph and the innovations described above, we generate cost functions to minimize the overall cost of traversing and the total time to reach a set of target locations.
VI-C1 Cost of Traversing
We formulate a cost of traversing with three main components: (a) edge cost, (b) vulnerability cost and (c) teaming rewards. Using the parameters from Table I, our edge cost is the positive valued $w_{e}$ for a team of robots to traverse edge $e$. For particularly vulnerable edges, it is desired for more agents to move together across the edge to enable moving in a formation for greater awareness. Thus, for each edge, to encode vulnerability, a minimum desired number of robots $a_{e}$ to traverse the edge is specified and an additional positive cost $m_{e}$ is incurred for each robot until the minimum is met. Finally, we consider an incentive for further teaming through positive rewards $r_{e}$ for additional robots moving together over the minimum desired. These components comprise the piecewise-linear cost of traversing an edge at a particular time, $C_{W_{e,t}}$, which can be be expressed as a function of the number of robots on the edge at that time, $p_{e,t}$. This cost function is depicted in Fig. 8(a).
$C_{W_{e,t}}=\begin{cases}0,&p_{e,t}=0\\ w_{e}+m_{e}(a_{e}-p_{e,t}),&0<p_{e,t}\leq a_{e}\\ w_{e}-r_{e}(p_{e,t}-a_{e}),&a_{e}\leq p_{e,t}\leq n_{A}\end{cases}$ (13)
(b) Cost of overwatch opportunity ${o}$ at time $t$ versus number of overwatch robots
Figure 8: Piecewise-Linear Cost Functions
@@ VI-C1 Cost of Traversing
To reduce computation time, we want a convex formulation of this cost function. The cost in (13) will not be convex in this case due to the zero point. To restate this cost, we first add an indicator variable for whether an edge is used, $\phi_{e,t}$, using Proposition 1. We then consider the function $f(p_{e,t})=w_{e}+m_{e}(a_{e}-p_{e,t})$, which is convex when we select $m_{e}\geq r_{e}$. Using Proposition 2, we can state the perspective function of $f(p_{e,t})$.
$\tilde{f}(p_{e,t},\phi_{e,t})=-m_{e}p_{e,t}+(w_{e}+m_{e}a_{e})\phi_{e,t}$ (14)
@@
By Proposition 2, the perspective function, $\tilde{f}(p_{e,t},\phi_{e,t})$, is a convex function in $p_{e,t}$ and $\phi_{e,t}$. Additionally, when $p_{e,t}=0$ then $\phi_{e,t}=0$ by definition, and this cost is zero. Otherwise, for any $p_{e,t}>0$ then $\phi_{e,t}=1$ and $\tilde{f}(p_{e,t})$ is identical to the second case in (13). Thus, this new function combines the first two cases in (13) and results in a convex piecewise linear function as follows.
$C_{W_{e,t}}=\begin{cases}-m_{e}p_{e,t}+(w_{e}+m_{e}a_{e})\phi_{e,t},&0\leq p_{e,t}\leq a_{e}\\ w_{e}-r_{e}(p_{e,t}-a_{e}),&a_{e}\leq p_{e,t}\leq n_{A}\end{cases}$ (15)
@@
Due to the convexity of this function, we can equivalently express this cost in our optimization problem with a linear term in the cost function that scales with our decision variables $\breve{C}_{W_{e,t}}$ and two linear constraints, as described in Proposition 3.
$\breve{C}_{W_{e,t}}$ $\geq-m_{e}p_{e,t}+(w_{e}+m_{e}a_{e})\phi_{e,t}$ (16)
$\breve{C}_{W_{e,t}}$ $\geq-r_{e}p_{e,t}+(w_{e}+r_{e}a_{e})\phi_{e,t}$ (17)
@@
These constraints are for the two line segments seen in Fig. 8(a). We utilize $\phi_{e,t}$ in (17), so that this term is defined whenever there are vehicles traversing and is zero otherwise. This sets the minimum value of the cost $\breve{C}_{W_{e,t}}$ to zero when vehicle’s are not traversing edge $e$ at time $t$. Since we are optimizing for minimum cost, $\breve{C}_{W_{e,t}}$ will be tight to the piecewise-linear cost function (15).
VI-C2 Cost of Overwatch
@@
When an edge can be monitored from a node, we consider this an overwatch opportunity and having overwatch while traversing can offer a cost reduction. When robots are traversing a corresponding edge, any number of robots at an overwatch node results in overwatch, but more robots may provide a greater reward. Using the parameters from Table I, we specify a positive $\omega_{o}$ as the benefit of overwatch opportunity ${o}$, $\alpha_{o}$ is the number of robots needed for full overwatch (i.e., to receive the full reward), and a positive $\gamma_{o}$ is a further reward for additional robots past $\alpha_{o}$ (i.e., a further teaming incentive). We consider $\rho_{{o},t}$ to be the number of overwatch robots for opportunity ${o}$ at time $t$. Thus, as depicted in Fig. 8(b), for a particular overwatch opportunity the “cost” is $C_{\Omega_{{o},t}}$.
$C_{\Omega_{{o},t}}=\begin{cases}-\frac{\omega_{o}}{\alpha_{o}}\rho_{{o},t},&0\leq\rho_{{o},t}\leq\alpha_{o}\\ -\omega_{o}-\gamma_{o}(\rho_{{o},t}-\alpha_{o}),&\alpha_{o}\leq\rho_{{o},t}\leq n_{A}\end{cases}$ (18)
@@
This “cost” will always be negative since it is rewarding overwatch. Similarly to the cost for traversing, since this piecewise-linear cost is convex when integrality constraints are relaxed and $\omega_{o}/\alpha_{o}\geq\gamma_{o}$, we can express the cost with a decision variable, $\breve{C}_{\Omega_{{o},t}}$, that enters linearly in our cost function with two linear constraints.
@@
$\breve{C}_{\Omega_{{o},t}}$ $\geq-\omega_{o}-\gamma_{o}(\rho_{{o},t}-\alpha_{o})$ (20)
We utilize the overwatch robot variables, $\rho_{{o},t}$, for ease of formulation, but can restate these constraints without these additional variables in order to reduce overall computation time for the optimization problem. For a particular ${o}\in{O}$, we consider the number of robots providing overwatch to be equal to the number of robots at the node, $v_{i}$, if there are robots on the edge, $e_{j}$. We set $\rho_{{o},t}=p_{v_{i},t}$, resulting in the following constraints.
@@
$\breve{C}_{\Omega_{{o},t}}$ $\geq-\omega_{o}-\gamma_{o}(p_{v_{i},t}-\alpha_{o})$ (22)
@@
$\breve{C}_{\Omega_{{o},t}}\geq-\frac{\omega_{o}}{\alpha_{o}}n_{A}p_{e_{j},t}$ (23)
@@
This constraint has a steeper slope (due to scaling by $n_{A}$) than (21) and (22) when there are robots on the edge (i.e., when $p_{e_{j},t}~{}>~{}0$), to ensure it is the least restrictive (which we can guarantee for convex (18)). When there are robots at the node and there are not robots on the edge (i.e., $p_{v_{i},t}~{}>~{}0$ and $p_{e_{j},t}~{}=~{}0$), (23) is the most restrictive overwatch constraint. This forces the cost $\breve{C}_{\Omega_{{o},t}}$ to be zero, since there is not a benefit from overwatch when there are not robots traversing the edge. Similarly, when $p_{v_{i}}=0$, $\breve{C}_{\Omega_{{o},t}}=0$ since there are not robots providing overwatch.
VI-C3 Cost of Time
@@
The final cost we consider, $C_{T_{t}}$, is for minimizing the time to achieve the goal. We formulate a cost that scales with time and is multiplied by the binary decision variables $\psi_{t}$ that represent whether robots have moved at time $t$. Thus, this cost is an incentive to achieve the goal as quickly as possible, since the total cost would be smallest when the robots accomplish the goal in minimal time. We define this cost for all time $t$.
$C_{T_{t}}=t\psi_{t}$ (24)
VI-C4 Overall Objective Function
@@
When we combine all of the cost terms, we express our overall objective function to minimize as follows.
$C=\sum_{t=1}^{n_{T}}\bigg{(}C_{T_{t}}+\sum_{e\in E}\breve{C}_{W_{e,t}}+\sum_{{o}\in{O}}\breve{C}_{\Omega_{{o},t}}\bigg{)}$ (25)
@@
A weight can be added to each term in (25) depending on the priority of a particular scenario (e.g., minimizing traversing cost versus minimizing time).
VI-D MIP Constraints
We add constraints to set support variables used in our cost functions and to restrict movement to the dynamic topological graph.
VI-D1 Edge Used Variables
@@
We utilize binary variables $\phi_{e,t}$ that track if an edge is used in Section VI-C1. We set $\phi_{e,t}$ to true if there are robots on edge $e$ at time $t$ and false otherwise through the following constraint.
$\phi_{e,t}\geq\frac{1}{n_{A}}p_{e,t}$ (26)
This constraint follows the form in Proposition 1 and $\phi_{e,t}$ is utilized in cases where cost is minimal when $\phi_{e,t}=0$.
VI-D2 Time Tracking Variables
@@
We add binary variables, $\psi_{t}$, to track if there are robots on the edges of the graph, to use in the time cost function in (24). Again we use the form in Proposition 1, but instead of for a single variable, we are considering the sum of the number of robots on edges at each time step. This sum tracks whether robots are still moving since robots cannot go instantaneously between nodes. Additionally, this formulation assumes that traversing edges has cost and remaining at nodes does not, such that remaining on an edge would not provide a benefit. Following Proposition 1, $\psi_{t}$ only contributes to increasing the overall cost in (24). For each time step $t$, we bound $\psi_{t}$.
$\psi_{t}\geq\frac{1}{n_{A}}\sum_{e\in E}p_{e,t}$ (27)
VI-D3 Start Locations

### v2

@@ V-B1 Cost of Traversing
$n_{E}$ Number of directional edges
@@
$n_{B}$ Number of start/begin locations
$n_{D}$ Number of goal/destination locations
@@
$B$ Set of start/begin locations $b$, $B\subseteq L$
@@
O{O} Set of overwatch opportunities o{o}, $(v_{i},e_{j})$ where node $v_{i}$ can overwatch edge $e_{j}$
@@
$n_{b}$ Number of robots at start location $b\in B$
@@
Cost of Traversing $\bar{w}_{e}$ Fixed cost to traverse edge $e\in E$
@@
$\alpha_{o}$ Number of robots for full overwatch for o{o}
$\gamma_{o}$ Reward for overwatch robots over $\alpha_{o}$ for o{o}
@@
$\breve{C}_{\bar{W}_{e,t}}$ Cont. 0 $\infty$ Cost of traversing edge $e$
$\breve{C}_{\Omega_{{o},t}}$ Cont. -$\infty$ 0 Cost of overwatch opportunity o{o}
@@
To reduce computation time, we want a convex formulation of this cost function. The cost in (6) is not convex due to the zero point. To restate this cost, we first add an indicator variable for whether an edge is used, $\phi_{e,t}$, using Proposition 1. We then consider the function $f(p_{e,t})=\bar{w}_{e}+m_{e}(a_{e}-p_{e,t})$, which is convex when we select $m_{e}\geq r_{e}$. Using Proposition 2, we can state the perspective function of $f(p_{e,t})$.
$\tilde{f}(p_{e,t},\phi_{e,t})=-m_{e}p_{e,t}+(\bar{w}_{e}+m_{e}a_{e})\phi_{e,t}.$ (7)
@@
By Proposition 2, the perspective function $\tilde{f}(p_{e,t},\phi_{e,t})$ is a convex function in $p_{e,t}$ and $\phi_{e,t}$. Additionally, when $p_{e,t}=0$ then $\phi_{e,t}=0$ by definition, and $\tilde{f}(p_{e,t},\phi_{e,t})=0$. Otherwise, for any $p_{e,t}>0$ then $\phi_{e,t}=1$ and $\tilde{f}(p_{e,t})$ is identical to the second case in (6). Thus, this new function combines the first two cases in (6) and results in a convex piecewise linear function as follows:
$C_{\bar{W}_{e,t}}=\begin{cases}-m_{e}p_{e,t}+(\bar{w}_{e}+m_{e}a_{e})\phi_{e,t},&0\leq p_{e,t}\leq a_{e}\\ \bar{w}_{e}-r_{e}(p_{e,t}-a_{e}),&a_{e}<p_{e,t}\leq n_{A}.\end{cases}$ (8)
@@
Due to the convexity of this function, we can equivalently express this cost in our optimization problem with a linear term in the cost function that scales with our decision variable $\breve{C}_{\bar{W}_{e,t}}$ and two linear constraints, as described in Proposition 3.
$\breve{C}_{\bar{W}_{e,t}}$ $\geq-m_{e}p_{e,t}+(\bar{w}_{e}+m_{e}a_{e})\phi_{e,t}$ (9)
$\breve{C}_{\bar{W}_{e,t}}$ $\geq-r_{e}p_{e,t}+(\bar{w}_{e}+r_{e}a_{e})\phi_{e,t}.$ (10)
@@
These constraints are for the two line segments seen in Fig. 3(a). We utilize $\phi_{e,t}$ in (10) as well, so that this term is defined whenever there are vehicles traversing and is zero otherwise. Overall, this ensures $\breve{C}_{\bar{W}_{e,t}}$ will go to zero when vehicles are not traversing edge $e$ at time $t$. Since we are optimizing for minimum cost, $\breve{C}_{\bar{W}_{e,t}}$ will be tight to the piecewise linear cost function (8).
V-B2 Cost of Overwatch
@@
When an edge can be monitored from a node, we consider this an overwatch opportunity. Having overwatch while traversing can offer a cost reduction proportional to the visibility of the edge from the overwatch node and the distance that can be monitored. When robots are traversing a corresponding edge, any number of robots at an overwatch node results in overwatch, but more robots may provide a greater reward. Using the parameters from Table I, we specify a positive $\omega_{o}$ as the benefit of overwatch opportunity o{o}, $\alpha_{o}$ is the number of robots needed for full overwatch (i.e., to receive the full reward), and a positive $\gamma_{o}$ is a further reward for additional robots past $\alpha_{o}$. We consider $\rho_{{o},t}$ to be the number of overwatch robots for opportunity o{o} at time $t$. Thus, as depicted in Fig. 3(b), for a particular overwatch opportunity the “cost” is $C_{\Omega_{{o},t}}$.
$C_{\Omega_{{o},t}}=\begin{cases}-\frac{\omega_{o}}{\alpha_{o}}\rho_{{o},t},&0\leq\rho_{{o},t}\leq\alpha_{o}\\ -\omega_{o}-\gamma_{o}(\rho_{{o},t}-\alpha_{o}),&\alpha_{o}<\rho_{{o},t}\leq n_{A}.\end{cases}$ (11)
@@
This “cost” will always be negative since it is rewarding overwatch. Similarly to the cost for traversing, since this piecewise linear cost is convex when integrality constraints are relaxed and $\omega_{o}/\alpha_{o}\geq\gamma_{o}$, we can express the cost with a decision variable, $\breve{C}_{\Omega_{{o},t}}$, that enters linearly in our cost function with two linear constraints.
@@
$\breve{C}_{\Omega_{{o},t}}$ $\geq-\omega_{o}-\gamma_{o}(\rho_{{o},t}-\alpha_{o}).$ (13)
We utilize the variables $\rho_{{o},t}$ for ease of formulation, but can restate these constraints without these additional variables to reduce computation time of the optimization problem. For a particular ${o}\in{O}$, we consider the number of robots providing overwatch to be equal to the number of robots at the node, $v_{i}$, if there are robots on the edge, $e_{j}$. We set $\rho_{{o},t}=p_{v_{i},t}$, resulting in the following constraints:
@@
$\breve{C}_{\Omega_{{o},t}}$ $\geq-\omega_{o}-\gamma_{o}(p_{v_{i},t}-\alpha_{o}).$ (15)
@@
$\breve{C}_{\Omega_{{o},t}}\geq-\frac{\omega_{o}}{\alpha_{o}}n_{A}p_{e_{j},t}.$ (16)
@@
This constraint has a steeper slope (due to scaling by $n_{A}$) than (14) and (15) when there are robots on the edge (i.e., when $p_{e_{j},t}~>~0$). This ensures it is the least restrictive (which we can guarantee for convex (11)). When there are robots at the node and there are not robots on the edge (i.e., $p_{v_{i},t}~>~0$ and $p_{e_{j},t}~=~0$), (16) is the most restrictive overwatch constraint. This forces the cost $\breve{C}_{\Omega_{{o},t}}$ to be zero, since there is not a benefit from overwatch when there are not robots traversing the edge. Similarly, when $p_{v_{i}}=0$, $\breve{C}_{\Omega_{{o},t}}=0$ since there are not robots providing overwatch.
V-B3 Cost of Time
@@
The final cost we consider, $C_{T_{t}}$, is for minimizing the time to achieve the goal. We formulate a cost that scales with time and is multiplied by the binary decision variables $\psi_{t}$ that represent whether robots have moved at time $t$. Thus, this cost, defined for all time $t$, is an incentive to achieve the goal as quickly as possible.
$C_{T_{t}}=t\psi_{t}.$ (17)
V-B4 Overall Objective Function
@@
When we combine all of the cost terms, we express our overall objective function to minimize as follows:
$C=\sum_{t=1}^{n_{T}}\bigg(C_{T_{t}}+\sum_{e\in E}\breve{C}_{\bar{W}_{e,t}}+\sum_{{o}\in{O}}\breve{C}_{\Omega_{{o},t}}\bigg).$ (18)
@@
Weights can be added to each term in (18) depending on the priority of a particular scenario (e.g., minimizing traversing cost versus minimizing time).
V-C MIP Constraints
We add constraints to set support variables used in our cost functions and to restrict movement to the DTG. These constraints assume that traversing an edge has cost and remaining at a node does not, such that remaining on an edge would not provide a benefit.
V-C1 Edge Used Variables
@@
We utilize binary variables $\phi_{e,t}$, that track if an edge is used, in cases where cost is minimal when $\phi_{e,t}=0$. Following Proposition 1, we set $\phi_{e,t}$ to $1$ if there are robots on edge $e$ at time $t$ and 0 otherwise.
$\phi_{e,t}\geq\frac{1}{n_{A}}p_{e,t}.$ (19)
V-C2 Time Tracking Variables
@@
We add binary variables, $\psi_{t}$, to track if there are robots on the edges of the graph at time $t$. We use the form in Proposition 1 with the sum of the number of robots on edges at each time step. This sum tracks whether robots are still moving since robots cannot go instantaneously between nodes. Following Proposition 1, $\psi_{t}$ only contributes to increasing cost. For each time step $t$, we bound $\psi_{t}$.
$\psi_{t}\geq\frac{1}{n_{A}}\sum_{e\in E}p_{e,t}.$ (20)
V-C3 Start Locations